# Long Zhao    Sanghyun Woo    Ziyu Wan    Yandong Li    Han Zhang    Boqing Gong    Hartwig Adam    Xuhui Jia    Ting Liu

### v3

@@ Abstract
In generative modeling, tokenization simplifies complex data into compact, structured representations, creating a more efficient, learnable space. For high-dimensional visual data, it reduces redundancy and emphasizes key features for high-quality generation. Current visual tokenization methods rely on a traditional autoencoder framework, where the encoder compresses data into latent representations, and the decoder reconstructs the original input. In this work, we offer a new perspective by proposing denoising as decoding, shifting from single-step reconstruction to iterative refinement. Specifically, we replace the decoder with a diffusion process that iteratively refines noise to recover the original image, guided by the latents provided by the encoder. We evaluate our approach by assessing both reconstruction (rFID) and generation quality (FID), comparing it to state-of-the-art autoencoding approaches. By adopting iterative reconstruction through diffusion, our autoencoder, namely $\epsilon$-VAE, achieves high reconstruction quality, which in turn enhances downstream generation quality by 22% and provides 2.3$\times$ inference speedup. We hope this work offers new insights into integrating iterative generation and autoencoding for improved compression and generation.
@@ 3.2 Objectives
Velocity prediction. We adopt the rectified flow parameterization, utilizing a linear optimization trajectory between data and noise, combined with velocity-matching objective:
$\mathbb{E}_{t\sim\pi(t),\bm{\epsilon}\sim\mathcal{N}(0,I)}\left[\|{\mathcal{G}}({\bm{x}}_{t},t,{\bm{z}})-(\bm{\epsilon}-{\bm{x}})\|^{2}\right].$ (9)
@@ 3.3 Noise and time scheduling
4 Experiments
@@
4.1 Reconstruction quality
@@
Resolution generalization. A notable feature of conventional autocencoders is their capacity to generalize and reconstruct images at higher resolutions during inference (Rombach et al., 2022). To assess this, we conduct inference on images with resolutions of $256\times 256$ and $512\times 512$, using $\epsilon$-VAE and VAE models trained at $128\times 128$. As shown in Tab. 1, $\epsilon$-VAE effectively generalizes to higher resolutions, consistently preserving its performance advantage over other VAEs. Furthermore, we find that fine-tuning models at the target (higher) resolution leads to improvement at it, which is consistent with the observation made by Sadat et al. (2024). We hence utilize this multi-stage training strategy in the following experiments when the target resolution is larger than $128\times 128$.
4.2 Class-conditional image generation
We now evaluate the generative performance of $\epsilon$-VAE when combined with latent diffusion models (Rombach et al., 2022). We perform standard class-conditional image generation tasks using the DiT-XL/2 model as our latent generative model (Peebles & Xie, 2023). Further details on the training setup are provided in Sec. C.3. Tab. 3 presents the image generation results of $\epsilon$-VAE and other competing VAEs at resolutions of $256\times 256$. The results show that $\epsilon$-VAE consistently outperforms other VAEs across different dowmsample factors. In addition, we emphasize that $\epsilon$-VAE achieves favorable generation quality while using only 25% of the token length typically required by SD-VAE. This token length reduction significantly accelerates latent diffusion model generation, leading to 2.3$\times$ higher inference throughput while maintaining competitive generation quality. These results confirm that the performance gains from the reconstruction task successfully transfer to the generation task, further validating the effectiveness of $\epsilon$-VAE.

### v4

# Epsilon-VAE: Denoising as Visual Decoding

@@ Abstract
In generative modeling, tokenization simplifies complex data into compact, structured representations, creating a more efficient, learnable space. For high-dimensional visual data, it reduces redundancy and emphasizes key features for high-quality generation. Current visual tokenization methods rely on a traditional autoencoder framework, where the encoder compresses data into latent representations, and the decoder reconstructs the original input. In this work, we offer a new perspective by proposing denoising as decoding, shifting from single-step reconstruction to iterative refinement. Specifically, we replace the decoder with a diffusion process that iteratively refines noise to recover the original image, guided by the latents provided by the encoder. We evaluate our approach by assessing both reconstruction (rFID) and generation quality (FID), comparing it to state-of-the-art autoencoding approaches. By adopting iterative reconstruction through diffusion, our autoencoder, namely $\epsilon$-VAE, achieves high reconstruction quality, which in turn enhances downstream generation quality by 22% at the same compression rates or provides 2.3$\times$ inference speedup through increasing compression rates. We hope this work offers new insights into integrating iterative generation and autoencoding for improved compression and generation.
@@ 3.2 Objectives
Velocity prediction. We adopt the rectified flow parameterization, utilizing a linear optimization trajectory between data and noise, combined with velocity-matching objective (Eq. 6). We inject the encoder output ${\bm{z}}$ into the objective by replacing $v_{\Theta}({\bm{x}}_{t},t)$ with ${\mathcal{G}}({\bm{x}}_{t},t,{\bm{z}})$.
@@
Note that applying LPIPS and adversarial losses on the estimated one-step sample could lead to potential objective bias. However, we would like to emphasize that $\epsilon$-VAE differs significantly from traditional diffusion models in that its diffusion decoder is conditioned on encoded latents ${\bm{z}}$. This conditioning provides a strong prior about the input image to reconstruct, resulting in a more accurate estimated one-step sample than in typical diffusion scenarios. Therefore, we believe the potential for objective bias is considerably reduced in $\epsilon$-VAE. Fine-tuning the diffusion decoder with frozen ${\bm{z}}$ like Sargent et al. (2025) could be a promising avenue for further improvement, which we will explore in our future work.
@@ 4.1 Reconstruction quality
Resolution generalization. A notable feature of conventional autoencoders is their capacity to generalize and reconstruct images at higher resolutions during inference (Rombach et al., 2022). To assess this, we conduct inference on images with resolutions of $256\times 256$ and $512\times 512$, using $\epsilon$-VAE and VAE models trained at $128\times 128$. As shown in Tab. 1, $\epsilon$-VAE effectively generalizes to higher resolutions, consistently preserving its performance advantage over other VAEs. Furthermore, we find that fine-tuning models at the target (higher) resolution leads to improvement at it, which is consistent with the observation made by Sadat et al. (2024). We hence utilize this multi-stage training strategy in the following experiments when the target resolution is larger than $128\times 128$.
@@
Table 3: Image reconstruction results of one-step $\epsilon$-VAE and SD-VAE on ImageNet $256\times 256$. A downsampling factor of $8\times 8$ is used for comparison. We include two variants of our model in the results: $\epsilon$-VAE-SD (B), which has a similar inference speed to SD-VAE, and $\epsilon$-VAE-SD (M), which matches SD-VAE in the number of parameters.
[…]
SD-VAE (Rombach et al., 2022) 0.74 25.68 0.820
$\epsilon$-VAE-SD (B) 0.57 25.91 0.826
$\epsilon$-VAE-SD (M) 0.51 26.45 0.830
One-step $\epsilon$-VAE. Note that the denoising process of $\epsilon$-VAE demonstrates promising results even with a single iteration. To show this, we provide a direct comparison between SD-VAE and our one-step $\epsilon$-VAE models in Tab. 3. This table presents image reconstruction quality on ImageNet $256\times 256$ with the $8\times 8$ downsampling factor. As shown, both $\epsilon$-VAE (B) and $\epsilon$-VAE (M) outperform SD-VAE across all metrics. These results confirm the effectiveness and efficiency of our one-step models compared to SD-VAE. Consequently, this allows $\epsilon$-VAE to be adapted for scenarios with latency-sensitive requirements, e.g., real-time visualization during image generation, by reducing the decoding step to a single pass.
4.2 Class-conditional image generation
@@
We now evaluate the generative performance of $\epsilon$-VAE when combined with latent diffusion models (Rombach et al., 2022). We perform standard class-conditional image generation tasks using the DiT-XL/2 model as our latent generative model (Peebles & Xie, 2023). Further details on the training setup are provided in Sec. C.3. Tab. 4 presents the image generation results of $\epsilon$-VAE and other competing VAEs at resolutions of $256\times 256$. The results show that $\epsilon$-VAE consistently outperforms other VAEs across different dowmsampling factors. In addition, we emphasize that $\epsilon$-VAE achieves favorable generation quality while using only 25% of the token length typically required by SD-VAE. This token length reduction significantly accelerates latent diffusion model generation, leading to 2.3$\times$ higher inference throughput while maintaining competitive generation quality. These results confirm that the performance gains from the reconstruction task successfully transfer to the generation task, further validating the effectiveness of $\epsilon$-VAE.
More importantly, $\epsilon$-VAE-SD achieves around 25% improvement in generation quality over SD-VAE at the $32\times 32$ downsampling factor, alongside a 3.2$\times$ inference speedup than SD-VAE at the $16\times 16$ downsampling factor with comparable FID. We observed similar training speedups for latent diffusion models utilizing $\epsilon$-VAE at this higher downsampling rate. These gains are more pronounced than those observed when increasing the downsampling factor from $8\times 8$ to $16\times 16$. These findings strongly suggest that the benefits of $\epsilon$-VAE and latent diffusion pipeline could be amplified with higher downsampling factors.
An additional advantage of scaling the autoencoder over the latent model lies in computational efficiency. Recent trends show latent diffusion models increasingly adopt Transformer architectures (Peebles & Xie, 2023), where self-attention scales quadratically with input resolution. In contrast, our convolution-based UNet decoder offers more favorable linear scaling. As models grow, shifting complexity to the autoencoder helps reduce the burden on the latent model, leading to a more efficient overall system.
@@ 4.3 Ablation studies
In Fig. 3 (right), reconstruction quality degrades when the number of denoising steps exceeds three. To enable large step sizes for the reverse process during inference, we introduce the denoising trajectory matching loss to implicitly model the conditional distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$, shifting the denoising distributions from traditional Gaussian to non-Gaussian multimodal forms (Xiao et al., 2022). However, the assumptions underlying this approach are most effective when the total number of denoising steps is small. This reveals an optimal range of one to three inference steps. The degradation beyond this range also suggests that uniform step spacing may no longer be ideal. Accordingly, we empirically explored alternative sampling strategies and found that a reversed logarithmic schedule yields improved performance, as shown in the figure.
@@ Appendix A Discussion
Multi-step vs. single-step decoding. While replacing single-step decoding with an iterative process may seem counter-intuitive due to increased computational cost, the diffusion-based decoder addresses this concern in three key ways. First, it offers scalable inference, where even a single-step variant already outperforms a plain VAE decoder, and additional steps further enhance quality (see Tab. 3). Second, it provides controllable trade-offs between computation and visual fidelity, allowing the number of steps to be adjusted at inference time based on application needs. Third, as shown in Tab. 4, it enables training under higher compression ratios, which helps offset the added cost of iterative decoding by reducing the size of latent representations.
@@ Appendix B Related work
While following a different motivation, Lee et al. (2023) essentially also proposes a VAE with a denoising decoder but uses the encoding as the “initial noise” instead of as conditioning for a standard diffusion model starting from a standard Gaussian distribution. This idea could be potentially used for speeding up the proposed approach, which we will explore in the future.
@@ Appendix D Additional experimental results
Results with classifier-free guidance. We provide additional results with classifier-free guidance (Ho & Salimans, 2022) under the $8\times 8$ downsample factor in Tab. 10. We find that $\epsilon$-VAE (M) performs relatively 20% better than SD-VAE and further improvements are obtained after we scale up our model to $\epsilon$-VAE (H). These results are consistent with the results without classifier-free guidance in Tab. 4, confirming the effectiveness of our model.
Table 10: Benchmarking class-conditional image generation on ImageNet $256\times 256$. We use the DiT-XL/2 architecture (Esser et al., 2024) for latent diffusion models and apply classifier-free guidance (Ho & Salimans, 2022).
[…]
$8\times 8$ SD-VAE (Rombach et al., 2022) 3.51
$\epsilon$-VAE-SD (M) 2.83
$\epsilon$-VAE-SD (H) 2.69